how many sequences of four heads and three tails can be obtained from seven coin tosses?
thank you very much.

Answer by stanbon(75887)

how many sequences of four heads and three tails can be obtained from seven coin tosses?
---------------
Just determine the number of ways to place the 4 heads; the tails will
take care of themselves the same number of ways.
----------------
Ans: 7C4 = 7*6*5*4/1*2*3*4 = 7C3 = 7*6*5/1*2*3 = 35
=======================================================
Cheers,
Stan H.